A metal rod of length 40.0cm at 20C is heated through a temperature of 45C.if the new length of the rod is 40.05cm calculate its linear expansivity


1
Expert's answer
2022-02-21T12:06:22-0500

Given:

"l_0=40.00\;\rm cm"

"\Delta T=25\:\rm K"

"l=40.05\;\rm cm"


The thermal linear expansion is described by an equation


"\Delta l=l_0\alpha\Delta T\\"


Hence, the linear expansivity of a metal

"\alpha=\frac{\Delta l}{l_0\Delta T}=\frac{0.05\:\rm cm}{40.00\:\rm cm*25\: K}=5*10^{-5}\: \rm K^{-1}"